# Neural Population Geometry and Optimal Coding of Tasks with Shared Latent Structure

### 26 Feb 2024 v1

@@ 5.1 Model of Multi-task Learning
We model a setting in which an agent learns a set of binary classification tasks by performing a linear readout on a set of neural population activities. Formally, we assume that each stimulus is associated with a $d$-dimensional latent variable $z_{\mu}\in\mathbb{R}^{d}$, with $1\leq\mu\leq p$ denoting the sample index. The labels for a specific task are formed by shattering the latent space using a hyperplane with a normal vector, $T$. Thus, we have that the binary classification task labels, $y_{\mu},$ satisfy the relation, $y_{\mu}=\mathrm{sign}(T\cdot z_{\mu})$, so that each $T$ vector defines a specific classification task. Associated with each latent, we also consider $n$-dimensional neural activity patterns, $x_{\mu}\in\mathbb{R}^{n}$. From these data, the agent then forms predictions of new data points using a supervised Hebbian readout rule [59, 32]. More precisely, given a new stimulus associated with latent variables $z_{+}$ and firing rates $x_{+}$, the agent forms a prediction $\hat{y}_{+}$ using the rule
@@
When the labels are balanced, this corresponds to using the difference in the mean activities for positively and negatively labeled examples (Fig. 1f). We evaluate how well the neural code can support downstream classification by calculating the generalization error of these predictions across different tasks (i.e., different $T$ vectors).
While we validate our theory on a wide range of data, we obtain our analytical results using a simplified Gaussian model. This is inspired by work on the Gaussian Equivalence Principle (GEP) in the deep learning theory community [35, 30, 31]. Formally, we assume that each pair of neural responses and latent variables, $(x,z)$ are jointly zero-mean Gaussians with covariance matrices:
@@ 5.2 Geometric Decomposition of Generalization Error Across Tasks
We prove a formula for the generalization error of the linear readout, averaged across different binary classification tasks (i.e., different $T$ vectors). To do so, we consider the limiting case in which the number of neurons, latent dimensions, and training samples are all large and of comparable size. Furthermore, we assume that the covariance matrices given above satisfy certain spectral properties–essentially we assume that their spectra decay at least as slowly as a power law with rate $\alpha<1$ (SM). Under these assumptions, we show that the average generalization error, $E_{g}$, is a decreasing function of the three geometric terms. Formally, we have
$E_{g}=\frac{1}{\pi}\tan^{-1}\bigg{(}\sqrt{\frac{\pi}{2pc^{2}\mathrm{PR}(\Psi)}+\frac{1}{ca}-1}\bigg{)},$ (3)
where we have introduced the total neural-latent correlation $c$, the alignment $a$, and the dimensionality of the population responses as measured by the participation ratio, $\mathrm{PR}(\Psi)$. These three terms correspond to statistics of the covariance matrices specified above:
$c$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})}{\mathrm{Tr}(\Psi)\mathrm{Tr}(\Omega)}$ (4)
$a$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})\mathrm{Tr}(\Psi)}{\mathrm{Tr}(\Phi^{T}\Psi\Phi)}$ (5)
@@
Since the function $F(w)=\frac{1}{\pi}\tan^{-1}\big{(}\sqrt{x-1}\big{)}$ is strictly increasing, we can see that each of these three geometric terms should be made as large as possible. Note that in writing Eq. (3), we have separated a term that depends on the number of training samples, $1/[c^{2}\operatorname{PR}(\Psi)]$ from a term which is independent of the sample size, $1/(ca)$. Thus we can see that the correlation and dimensionality terms become less important as $p$ grows.
@@ 5.4 Optimal Codes
We derive which neuronal codes achieve the lowest generalization error, given a fixed number of samples to train on and a fixed latent structure. Specifically, we calculate which neuron-neuron and neuron-latent covariance matrices, $\Psi$, $\Phi$ achieve the lowest multi-task generalization error, given fixed a fixed latent covariance $\Omega$ and training set size $p$ (SM). We do this by optimizing Eq (3) with respect to $\Psi$ and $\Phi$ subject to the constraint that the entire covariance matrix between neurons and latents be positive semi-definite, a necessary condition for the code to be realizable. Using this approach, we find that the left and right singular vectors of $\Phi$ are the eigenvectors of $\Psi$ and $\Omega$, respectively for the optimal code. This shows that independent directions in the latent space map directly onto independent directions in the firing rate space. Furthermore, we obtain the following simple formula for the eigenvalues of the optimal neural code. Denoting the eigenvalues of $\Psi$ and $\Omega$ as $\psi_{i}$ and $\omega_{i}$ respectively, we have up to a permutation symmetry:
$\psi_{i}=C\frac{\omega_{i}}{2p\omega_{i}+\sum_{k}\omega_{k}},$ (7)
@@ 5.5 MLP Experiments
We used random and trained MLPs to test several predictions from our theory using explicitly non-Gaussian artificial neural response data. To generate these data, we first sampled a set of $d=40$ dimensional latent variables from a multivariate Gaussian distribution with eigenvalues $\omega_{k}=k^{-0.2}$, though we also show similar results for a flat spectrum (SM). For these experiments, we sampled $5\cdot 10^{5}$ latent variables independently from this distribution as a training dataset. Using these latent variables we generated a set of $N_{t}=500$ tasks by randomly shattering the latent space. Latent variables were then passed through a $3$ layer perceptron with randomly initialized weights. We chose the size of each intermediate layer to be twice the size of the previous one to ensure the dimension of the representation did not decrease. Each layer was composed of a linear transform, followed by a relu non-linearity (see SM for results using a tanh non-linearity). After sampling the task labels and passing the latents through the random MLP, we trained a $4$ hidden-layer MLP to predict task labels from the random MLP responses. The trained network was similarly composed of linear-batchnorm-relu blocks, and we used the adam optimizer to train the network [60] through a single epoch. This setup corresponds to a multi-task version of the hidden manifold modeling framework studied in the machine learning community [30].
@@ 5.6 Macaque Analyses
We drew from a publicly available dataset containing multi-unit recordings from V4 and IT taken from 2 monkeys [33]. These recordings were taken as the monkeys viewed visual stimuli as described in [33] and contained 88 V4 and 168 IT neural sites. The stimuli for this task were drawn from a generative model in which a total of 64 objects coming from 8 categories were displayed against varying backgrounds. These images were generated by varying $d=6$ continuous latent variables that controlled the object size, angle, and position. Latent variables were drawn iid from a mean zero uniform distribution.
We tested the linear decodability of these latent variables from the neural firing rates using the scheme described in Sec. 5.1. For the raw pixels, we first carried out a Gaussian random projection onto a $5,000$ dimensional space before applying this scheme. Specifically, for each task condition and for each of the 8 object category types, we formed task labels for $N_{t}=300$ tasks by randomly shattering the latent space. For each condition and category type, there were $320$ examples per condition and object category type. We z-scored the latent variables to ensure that there was no especially informative latent variable. We then formed predictions using the supervised Hebbian readout described in Sec. 5.1 after mean centering the neuronal firing rates. This procedure was repeated over 15 different train/test data splits. To generate the results in Fig. 7, we then averaged the generalization error across the $N_{t}$ classification tasks, data splits, conditions, and image categories. Error bars denote the standard error of the mean for the estimated generalization error across the $c=8$ category types. In the SM, we also show the geometry and generalization errors for each of the categories individually, as well as the error obtained by pooling all categories together.
@@ Acknowledgements
We thank Thomas Yerxa for comments on this manuscript. This work was funded by the Center for Computational Neuroscience at the Flatiron Institute of the Simons Foundation and the Klingenstein-Simons Award to S.C. All experiments were performed on the Flatiron Institute high-performance computing cluster.

### 11 Apr 2024 v2

@@ 5.1 Model of multi-task learning
We model a setting in which an agent learns a set of binary classification tasks by performing a linear readout on a set of neural population activities. Formally, we assume that each stimulus is associated with a $d$-dimensional latent variable $z_{\mu}\in\mathbb{R}^{d}$, with $1\leq\mu\leq p$ denoting the sample index. The labels for a specific task are formed by shattering the latent space using a hyperplane with a normal vector, $T$. Thus, we have that the binary classification task labels, $y_{\mu},$ satisfy the relation, $y_{\mu}=\mathrm{sign}(T\cdot z_{\mu})$, so that each $T$ vector defines a specific classification task. Associated with each latent, we also consider $n$-dimensional neural activity patterns, $x_{\mu}\in\mathbb{R}^{n}$. From these data, the agent then forms predictions of new data points using a supervised Hebbian readout rule [38]. More precisely, given a new stimulus associated with latent variables $z_{+}$ and firing rates $x_{+}$, the agent forms a prediction $\hat{y}_{+}$ using the rule
@@
When the labels are balanced, this corresponds to using the difference in the mean activities for positively and negatively labeled examples (Fig. 1f). We evaluate how well the neural code can support downstream classification by calculating the generalization error of these predictions, averaged across different tasks (i.e., different $T$ vectors), though we also calculate the error for a fixed $T$ along the way (SM).
While we validate our theory on a wide range of data, we obtain our analytical results using a simplified Gaussian model. This is inspired by work on the Gaussian Equivalence Principle (GEP) in deep learning theory [40, 36, 37]. Formally, we assume that each pair of neural responses and latent variables, $(x,z)$ are jointly zero-mean Gaussians with covariance matrices:
@@ 5.2 Geometric decomposition of generalization error across tasks
We prove a formula for the generalization error of the linear readout, averaged across different binary classification tasks (i.e., different $T$ vectors). To do so, we consider the limiting case in which the number of neurons, latent dimensions, and training samples are all large and of comparable size. Furthermore, we assume that the covariance matrices given above satisfy certain spectral properties (SM). Under these assumptions, we show that the average generalization error, $E_{g}$, is a decreasing function of the four geometric terms. Formally, we have
$E_{g}=\frac{1}{\pi}\tan^{-1}\bigg{(}\sqrt{\frac{\pi}{2pc^{2}\mathrm{PR}(\Psi)}+\frac{1}{f}+\frac{1}{s}-1}\bigg{)},$ (3)
where we have introduced the total neural-latent correlation $c$, the signal-signal factorization $f$, the signal-noise factorization $s$, and the dimension of the population responses as measured by the participation ratio, $\mathrm{PR}(\Psi)$. These terms correspond to statistics of the covariance matrices specified above:
$c$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})}{\mathrm{Tr}(\Psi)\mathrm{Tr}(\Omega)}$ (4)
@@
$f$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})^{2}}{\mathrm{Tr}(\Omega)\mathrm{Tr}(\Phi^{\top}\Phi\Omega^{-1}\Phi^{\top}\Phi)}$ (6)
$s$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})^{2}}{\mathrm{Tr}(\Omega)\mathrm{Tr}(\Phi^{\top}(\Psi-\Phi\Omega^{-1}\Phi^{\top})\Phi)}$ (7)
Since the function $F(w)=\frac{1}{\pi}\tan^{-1}\big{(}\sqrt{w-1}\big{)}$ is strictly increasing, we can see that each of these geometric terms should be made as large as possible. As discussed in the SM, the term $f$ measures the overall degree of orthogonality between coding directions, while $s$ measures the amount of noise that lies along the signal directions. The neural noise is described by the noise covariance matrix, $\mathrm{cov}(x|z)=\Psi-\Phi\Omega^{-1}\Phi^{\top},$ which appears in the definition above. Importantly, this neural noise may include variability that is related to latent variables that are not measured experimentally. Motivated by this observation, we describe how these two factorization terms can be collapsed into a single factorization term in the SM. Note that in writing Eq. (3), we have separated a term that depends on the number of training samples, $1/[c^{2}\operatorname{PR}(\Psi)]$ from a term which is independent of the sample size, $1/f+1/s$. Thus we can see that the correlation and dimension terms become less important as $p$ grows.
@@ 5.4 Optimal codes
We derive which neuronal codes achieve the lowest generalization error, given a fixed number of samples to train on and a fixed latent structure. Specifically, we calculate which neuron-neuron and neuron-latent covariance matrices, $\Psi$, $\Phi$ achieve the lowest multi-task generalization error, given fixed a fixed latent covariance $\Omega$ and training set size $p$ (SM). We do this by optimizing Eq (3) with respect to $\Psi$ and $\Phi$ subject to the constraint that the entire covariance matrix between neurons and latents be positive semi-definite, a necessary condition for the code to be realizable. Using this approach, we find that the left and right singular vectors of $\Phi$ are the eigenvectors of $\Psi$ and $\Omega$, respectively for the optimal code. This shows that independent directions in the latent space map directly onto independent directions in the firing rate space (SM). Furthermore, we obtain the following simple formula for the eigenvalues of the optimal neural code. Denoting the eigenvalues of $\Psi$ and $\Omega$ as $\psi_{i}$ and $\omega_{i}$ respectively, we have up to a permutation symmetry:
$\psi_{i}=C\frac{\omega_{i}}{2p\omega_{i}+\pi\sum_{k}\omega_{k}},$ (8)
@@ 5.5 MLP Experiments
We used random and trained MLPs to test several predictions from our theory using explicitly non-Gaussian artificial neural response data. To generate these data, we first sampled a set of $d=40$ dimensional latent variables from a multivariate Gaussian distribution with eigenvalues $\omega_{k}=k^{-0.2}$. For these experiments, we sampled $5\cdot 10^{5}$ latent variables independently from this distribution as a training dataset. Using these latent variables we generated a set of $N_{t}=500$ tasks by randomly shattering the latent space. Latent variables were then passed through a $3$ layer perceptron with randomly initialized weights. We chose the size of each intermediate layer to be twice the size of the previous one to ensure the dimension of the representation did not decrease. Each layer was composed of a linear transform, followed by a relu non-linearity (see SM Figs. 2&3 for results using a tanh non-linearity). After sampling the task labels and passing the latents through the random MLP, we trained a $4$ hidden-layer MLP to predict task labels from the random MLP responses. The trained network was similarly composed of linear-batchnorm-relu blocks, and we used the adam optimizer to train the network [73] through a single epoch. This setup corresponds to a multi-task version of the hidden manifold modeling framework studied in deep learning theory [36].
@@ 5.6 Macaque analyses
We drew from a publicly available dataset containing multi-unit recordings from V4 and IT taken from 2 monkeys [39]. These recordings were taken as the monkeys viewed visual stimuli as described in [39] and contained 88 V4 and 168 IT neural sites, though we reproduce all results projecting IT and pixel responses down to 88 randomly chosen dimensions (SM Fig. 4). The stimuli for this task were drawn from a generative model in which a total of 64 objects coming from 8 categories were displayed against varying backgrounds. These images were generated by varying $d=6$ continuous latent variables that controlled the object size, angle, and position. Latent variables were drawn iid from a mean zero uniform distribution.
We tested the linear decodability of these latent variables from the neural firing rates using the scheme described in Sec. 5.1. For the raw pixels, we first carried out a Gaussian random projection onto a $5,000$ dimensional space before applying this scheme. Specifically, for each task condition and for each of the 8 object category types, we formed task labels for $N_{t}=300$ tasks by randomly shattering the latent space. For each condition and category type, there were $320$ examples per condition and object category type. We z-scored the latent variables to ensure that there was no especially informative latent variable. We then formed predictions using the supervised Hebbian readout described in Sec. 5.1 after mean centering the neuronal firing rates. This procedure was repeated over 15 different train/test data splits. To generate the results in Fig. 7, we then averaged the generalization error across the $N_{t}$ classification tasks, data splits, conditions, and image categories. Error bars denote the standard error of the mean for the estimated generalization error across the $c=8$ category types. In SM, we also show the geometry and generalization errors for each of the categories individually (SM Figs. 5&6), as well as the error obtained by pooling all categories together (SM Fig. 7).
@@ Acknowledgements
We thank Thomas Yerxa, Chi-Ning Chou, Arnav Raha, and Jenelle Feather for helpful discussions and comments on an earlier version of this manuscript. This work was funded by the Center for Computational Neuroscience at the Flatiron Institute of the Simons Foundation. S.C. is also partially supported by a Sloan Research Fellowship, and a Klingenstein-Simons Award. All experiments were performed on the high-performance computing cluster at the Flatiron Institute.